# On the extraction of Alcock-Paczyński signal from voids: a novel approach via reconstruction

### v1

@@ 2.5 Cosmological reconstruction
The reconstruction is performed in several steps. In the first step, the positions of galaxies are interpolated onto a cubic grid to estimate their number density, $\rho_{\mathrm{obs}}$. The cube is large enough to encompass the survey volume. The procedure is repeated to estimate, at the same positions, the number density of the random objects $\rho_{\mathrm{r}}$. The overdensity field is then computed as: $\delta_{\mathrm{obs}}=\frac{\rho_{\mathrm{obs}}}{\rho_{\mathrm{r}}}-1$. Since a linear bias is assumed, the mass overdensity can be estimated by simply rescaling the galaxy overdensity by the bias factor $b$. In cells where $\rho_{\mathrm{r}}=0$, the overdensity is set to zero to avoid numerical instabilities. In the second step, a Gaussian filter is applied to the overdensity field. Since this operation corresponds to a simple multiplication in Fourier space, we first Fourier transform the overdensity field and multiply it by the Gaussian filter $\mathcal{S}(\mathbf{k})=e^{-k^{2}R_{\mathrm{s}}^{2}/2}$, where $R_{\mathrm{s}}$ is the smoothing scale. Finally, we inverse Fourier transform the smoothed field back to configuration space. In the third, and central step, it evaluates the displacement field $\boldsymbol{\psi}_{\mathrm{RSD}}$ by solving the system of equations (Nusser 1994):
@@ 3 Data
3.1 The simulated halo catalogues
@@
To match the number density of the LOWZ and CMASS samples in SDSS-III DR12 (Reid et al. 2016), we apply a mass cut of $m_{\mathrm{min}}=10^{13}\ h^{-1}M_{\odot}$, yielding about $4\times 10^{5}$ halos per catalog.
We also compute the fiducial growth rate at $z_{\mathrm{snap}}$ as $f=\Omega_{\mathrm{m}}^{0.55}=0.763$, and determine the effective halo bias to be $b=1.87\pm 0.03$, from their measured 2-point correlation function, as detailed in Appendix C. These values yield a distortion parameter $\beta=f/b=0.41$, which is used in the reconstruction process (Section 2.5).
To build the three types of catalogs used in our analysis, we take the positions of the halo in real space and move them along the LOS to their redshift-space position using Eq. 9 combined with the peculiar velocities provided by the simulation. We then apply the reconstruction procedure to the redshift-space halo catalog (see Section 3.2). As a result, we obtain three distinct types of halo catalogs: one in real space, one in redshift space, and one in reconstructed space. We then apply the void finder to each of these catalogs (see Section 3.4), producing the three corresponding void catalogs used in our cross-correlation analyses.
@@ 3.2 The redshift space halo catalogs
To generate the redshift-space catalogs from the real-space data described in Section 3.1, we adopt the distant observer approximation, aligning the LOS with the Cartesian $z$-axis. The center of the box is placed at a comoving distance $d(z_{\mathrm{snap}})$, as given by Eq. (1), and distances to all halos are computed accordingly. The observed redshift of each halo is then calculated using Eq. (8), combining the cosmological redshift with the Doppler shift from the LOS component of peculiar velocities, $z_{d}=\mathbf{v}_{\parallel}/c$, provided by the simulation. We use Eq. 1 to estimate the new comoving distances and their Cartesian coordinates $(x,y,z)$. Since this transformation can shift some halos beyond the cube boundaries, we trim the catalog to exclude those objects. The resulting loss is negligible, affecting only $\sim 0.3\%$ of halos.
These redshift-space mocks are used both for the redshift space analysis and as input to the reconstruction algorithm to extract the reconstructed-space catalogs.
@@ 3.3 The reconstructed halo catalogs
Reconstructed space halo catalogs were obtained by applying the Zel’dovich reconstruction algorithm to the redshift-space halo catalogs. The reconstruction procedure (Section 2.5) requires defining a cubic grid of cells to interpolate the halo number density. We set $N_{\mathrm{cell}}=256$, corresponding to cell size of $3.9\ h^{-1}\mathrm{Mpc}$. The density field is then smoothed with a Gaussian filter of radius $R_{\mathrm{s}}\sim 5\ h^{-1}\mathrm{Mpc}$ to minimizes the residual RSD effects. The choice of $R_{\mathrm{s}}$ is motivated in the Appendix A. As for the number of cells, we verified that its choice has no significant impact on the reconstruction as long as the size of the cell is smaller than the smoothing radius (White 2021). The fact that the smoothing radius is just 1.5 smaller than that of the smallest void in the sample ($\sim 7.5\ h^{-1}\mathrm{Mpc}$) has no significant impact on our results. The role of the smoothing is to minimize the effects of nonlinearities when removing RSDs, which it does effectively, as shown in the following sections. Importantly, it plays no role in the void detection process, which relies on a Voronoi tessellation of the discrete tracer distribution. As in the previous section, we remove all halos displaced outside the cube resulting in a loss of $\sim 0.4\%$ with respect to the original catalog in real space.
@@ 3.4 The void catalogs
Void analyses identify these structures in a 3-dimensional distribution of mass tracers, in our case simulated dark matter halos. Several void identification algorithms have been proposed in the literature, (see e.g. Colberg et al. 2008; Cautun et al. 2014, for a cross-comparison of various techniques available at that time). In this work, we identify voids in the simulated halo catalogs with the publicly available code VIDE  [^4] https://bitbucket.org/cosmicvoids/vide_public/ (Sutter et al. 2015), based on the watershed algorithm ZOBOV (Neyrinck 2008), which has been widely tested and applied in the literature (e.g. Hamaus et al. 2020; Contarini et al. 2023; Verza et al. 2024). This void finder estimates local densities using Voronoi tessellation and identifies voids as basins in the tracer field. The output is a catalog of voids, each described by its center,
@@
calculated as the volume-weighted barycenters, and the effective radius,
@@
We apply VIDE to all three types of catalogs (real, redshift, and reconstructed space) across 100 mocks. Fig. 1 on the top panel shows the void size function, i.e., the abundance of voids as a function of their effective radius calculated as the number density of voids $n_{\mathrm{v}}$ per unit radius interval, in real (green stars), redshift (orange triangles), and reconstructed (blue dots) space. The solid lines represent the mean across 100 mocks, while the shaded regions indicate the standard deviation. The results in real and reconstructed space are in good agreement within errors, whereas redshift space shows a slight mismatch, as can be better appreciated by looking at the difference with the real space void size function, normalized by the $1\sigma$ error, represented in the bottom panel of Fig. 1. The discrepancies between redshift and real space are particularly prominent at intermediate scales $20,40]\ h^{-1}$Mpc where the voids are more abundant, while residuals between real and reconstructed space do not show a dependence on the size, always being within the $1\sigma$ interval (grey band). This suggests that the reconstruction procedure successfully recovers the statistical properties of the void distribution, including both their abundance and size, while redshift space distortions affect their identification. A possible explanation is that voids squished along the LOS in real space are more likely to not being selected in real space while identified in redshift space, as discussed in Correa et al. (2022). Additionally, void finders identify spurious, shallow voids due to shot noise from the discrete tracer distribution, commonly referred to as Poisson voids (Neyrinck 2008; Cousinou et al. 2019). These will impact void statistics such as the void size function and the VGCF. A standard mitigation is to exclude small voids below a fixed multiple of the mean tracer separation (mps) as in Hamaus et al. (2020), with values typically around 3 times the mps. However, such cuts are somewhat arbitrary and not fully effective at eliminating contaminants, while considerably reducing the sample size (Cousinou et al. 2019). For this reason, we choose to retain all identified voids and test the robustness of our analysis by comparing results with and without small voids (see Section 5.2). A more detailed investigation into the impact of spurious voids is left for future work.
@@ 4.1 Estimator
The multipoles of the VGCF are obtained using the Davis-Peebles (DP) auto-correlation function estimator (Davis & Peebles 1983).
The DP estimator is based on counting object pairs at a given separation and compare them with pair counts performed in a population of synthetic objects distributed over the same volume and sharing the same selection effects as the real ones but with no intrinsic spatial correlation properties, the so-called random sample.
@@
where $\mathcal{D}_{\text{v}}\mathcal{D}_{\mathrm{h}}$ are the void-halo (or more generally void-tracer) pair counts at radial $r$ and angular $\mu$ separation, and $\mathcal{D}_{\text{v}}\mathcal{R}_{\mathrm{h}}$ are the void-random pair counts. The quantities $n_{\mathrm{R}}$ and $n_{\mathrm{h}}$ represent the total number counts of the random objects and of halos respectively. In Eq. (22) $r$ is a dimensionless quantity representing the physical separation between the center of the void and the halo divided by the void effective radius $R$. In this analysis the pair counts are evaluated in 25 equally-spaced bins in the range $[0,3]$ while the cosine angle $\mu$ is evaluated in $100$ bins between $[0,1]$. In our analysis, we compute the cross pair with the random catalogue by generating a single catalog of synthetic objects uniformly distributed across the volume, with a number density 20 times higher than that of the halos. This random catalog is used for all cross-correlation analyses, as they share the same geometry and mean halo density.
Once the VGCF is evaluated, the multipoles can be computed using Eq. (16). Particular attention is given to the quadrupole, as it serves as a crucial diagnostic tool for understanding the effects of RSD. In real space, where no RSD are present, the quadrupole signal is expected to vanish due to the spherical symmetry of the stacked void profile. Thus, a zero quadrupole signal is used as a benchmark to assess the accuracy of the reconstruction procedure within the appropriate cosmological model. On top of the impact of RSD on the quadrupole shape, AP geometric distortions cause the quadrupole to deviate from zero. The direction of the deviation, compression or elongation, depends on whether the parameter $\varepsilon$, defined by Eq. (4), is lower or greater than 1. As we use the fiducial cosmology of the simulation in the distance estimate, no genuine AP-induced distortion is expected, and $\varepsilon$ should therefore be unity.
@@ 4.2 Likelihood
To sample the posterior probability distribution of all model parameters we make use of the affine-invariant Markov chain Monte Carlo (MCMC) ensemble sampler emcee (Foreman-Mackey et al. 2019). The quality of the maximum-likelihood model (best fit) is assessed via evaluation of the reduced $\chi^{2}$ statistics:
@@ 4.3 Covariance matrix
All analyses are carried out using a dedicated covariance matrix for each mock, estimated via jackknife resampling of its corresponding void catalog. Given that the number of available mocks is insufficient to reliably compute a full covariance matrix across realizations, we follow a standard approach adopted in the literature and apply a delete-1 jackknife strategy to estimate the covariance for each of the 100 mocks individually (see e.g. Hamaus et al. 2020, 2022; Radinović et al. 2023). In Appendix A.2, we show that the jackknife-derived covariance matrices are consistent with those obtained from the full mock-to-mock covariance. Based on this agreement, we adopt the jackknife covariance, $\text{Cov}[\boldsymbol{\xi}]_{ij}$, throughout our analysis.
This methodology relies on ergodicity, which allows us to average measurements across different spatial patches to estimate the covariance matrix (Hamaus et al. 2020). In this approach, we remove one void at a time from the sample when estimating $\xi_{\ell}(r)$, providing a total of $N_{\mathrm{v}}$ jackknife samples, where $N_{\mathrm{v}}$ is the total number of voids in the catalog. These samples can then be used in Eq. (27) to calculate $\text{Cov}[\boldsymbol{\xi}]_{ij}$, with an additional normalization factor of $(N_{\mathrm{v}}-1)$ to account for the statistical weight of the jackknife sample size. The elements of the covariance matrix are computed as follows:
@@ 5 Results
  • The first, a pre-requisite for the second, is to assess the performance of the reconstruction algorithm in removing RSDs. To do so, we focus on the parameter $\beta$, which quantifies the amplitude of RSD effects in the VGCF multipoles. In the ideal case of perfect RSD removal, we expect $\beta=0$. In practice, however, uncertainties are expected to arise from imperfect dynamical reconstruction, void detection, and VGCF estimation. To isolate the effects introduced solely by reconstruction and assess their impact on the void-finding procedure, we compare the best-fit values of $\beta$ obtained from the reconstructed catalogs with those measured directly in real space, rather than with the expected value $\beta=0$. A close match between the two $\beta$ values would indicate that the reconstruction effectively removes RSDs.
  • The second is to evaluate whether performing the AP test in reconstructed space yields tighter constraints on the geometric distortion parameter $\varepsilon$ compared to using voids identified in redshift space. To this end, we compare the AP constraints derived from both redshift and reconstructed spaces void catalogs using the full void sample (Section 5.1). We then assess the impact of excluding small voids, motivated by their greater susceptibility to noise and systematics, by repeating the analysis on size-selected subsamples in both spaces (Section 5.2).
Finally, to assess the robustness of the results presented in this section, we perform a series of consistency tests. These include examining the sensitivity of the method to the parameters used in the reconstruction process and evaluating its stability with respect to the choice of the covariance matrix. A detailed description of these tests and their outcomes is provided in Appendix A.
@@
Fig. 2, shows the three measured multipoles averaged for the 100 mocks (described in Section 3), in redshift space (orange triangles and solid curves), reconstructed space (blue dots and solid curves) and real space (green stars and solid curves). Focusing on the quadrupole (central panel), we observe a positive bump near to the ridge at $r\sim 1$ in redshift space. This feature signals the presence of coherent outflows and manifests as an apparent elongation of the void along the LOS. On the other hand, the quadrupole measured in reconstructed space is compatible within error bars with the one in real space, with both being consistent with zero, indicating the effective removal of RSDs. The consistency of the hexadecapole with zero further supports this conclusion.
@@
One of the central motivations of this work is to develop a methodology that enables the estimation of the geometric distortion parameter $\varepsilon$ independently of the systematics that affect the measurement of $\beta$ in redshift space, systematics that can drive $\beta$ far from its true value. Disentangling these contributions is particularly important because, in redshift space, the parameters $\varepsilon$ and $\beta$ are strongly correlated. This is evident in Fig. 3, where their joint posterior displays a pronounced alignment, with a correlation coefficient of $\mathcal{R}=0.67$ estimated directly from the parameter chains. Since performing a robust AP test requires an accurate treatment of RSD, applying reconstruction methods proves highly effective: as shown in Fig. 4, the posterior of $\varepsilon$ and $\beta$ becomes nearly spherical after reconstruction, and the correlation is reduced to $\mathcal{R}=0.23$, demonstrating that the degeneracy between RSD and AP effects is largely mitigated.
@@
Table 1: Table summarizing the values of the parameters $\varepsilon$ and $\beta$, associated to their estimated errors, for the analyses in redshift, reconstructed, and real space. The precision in reconstructed space is increased by $\sim 23\%$ with respect to redshift space, and approaches the real-space analysis precision.

### v2

@@ 2.5 Cosmological reconstruction
The reconstruction is performed in several steps. In the first step, the positions of galaxies are interpolated onto a cubic grid to estimate their number density, $\rho_{\mathrm{obs}}$. The cube is large enough to encompass the survey volume. The procedure is repeated to estimate, at the same positions, the number density of the random objects $\rho_{\mathrm{r}}$. The overdensity field is then computed as: $\delta_{\mathrm{obs}}=\frac{\rho_{\mathrm{obs}}}{\rho_{\mathrm{r}}}-1$. Since a linear bias is assumed, the mass overdensity can be estimated by simply rescaling the galaxy overdensity by the bias factor $b$. In cells where $\rho_{\mathrm{r}}=0$, the overdensity is set to zero to avoid numerical instabilities. In the second step, a Gaussian filter is applied to the overdensity field. Since this operation corresponds to a simple multiplication in Fourier space, we first Fourier transform the overdensity field and multiply it by the Gaussian filter $\mathcal{S}(\mathbf{k})=e^{-k^{2}R_{\mathrm{s}}^{2}/2}$, where $R_{\mathrm{s}}$ is the smoothing scale. Finally, we inverse Fourier transform the smoothed field back to configuration space. In the third, and central, step it evaluates the displacement field $\boldsymbol{\psi}_{\mathrm{RSD}}$ by solving the system of equations (Nusser 1994):
@@ 3 Data
3.1 The simulated halo catalogs
@@
We also computed the fiducial growth rate at $z_{\mathrm{snap}}$ as $f=\Omega_{\mathrm{m}}^{0.55}=0.763$, and determined the effective halo bias to be $b=1.87\pm 0.03$, from the measured 2-point correlation function, as detailed in Appendix C. These values yield a distortion parameter $\beta=f/b=0.41$, which is used in the reconstruction process (Section 2.5).
To build the three types of catalogs used in our analysis, we took the positions of the halo in real space and move them along the LOS to their redshift-space position using Eq. 9 combined with the peculiar velocities provided by the simulation. We then applied the reconstruction procedure to the redshift-space halo catalog (see Section 3.2). As a result, we obtained three distinct types of halo catalogs: one in real space, one in redshift space, and one in reconstructed space. We then apply the void finder to each of these catalogs (see Section 3.4), producing the three corresponding void catalogs used in our cross-correlation analyses.
@@ 3.2 The redshift space halo catalogs
To generate the redshift-space catalogs from the real-space data described in Section 3.1, we adopted the distant observer approximation, aligning the LOS with the Cartesian $z$-axis. The center of the box is placed at a comoving distance $d(z_{\mathrm{snap}})$, as given by Eq. (1), and distances to all halos are computed accordingly. The observed redshift of each halo is then calculated using Eq. (8), combining the cosmological redshift with the Doppler shift from the LOS component of peculiar velocities, $z_{d}=\mathbf{v}_{\parallel}/c$, provided by the simulation. We used Eq. 1 to estimate the new comoving distances and their Cartesian coordinates $(x,y,z)$. Since this transformation can shift some halos beyond the cube boundaries, we trimmed the catalog to exclude those objects. The resulting loss is negligible, affecting only $\sim 0.3\%$ of halos. These redshift-space mocks are used both for the redshift space analysis and as input to the reconstruction algorithm to extract the reconstructed-space catalogs.
@@ 3.3 The reconstructed halo catalogs
Reconstructed space halo catalogs were obtained by applying the Zel’dovich reconstruction algorithm to the redshift-space halo catalogs. The reconstruction procedure (Section 2.5) requires defining a cubic grid of cells to interpolate the halo number density. We set $N_{\mathrm{cell}}=256$, corresponding to cell size of $3.9\ h^{-1}\mathrm{Mpc}$. The density field is then smoothed with a Gaussian filter of radius $R_{\mathrm{s}}\sim 5\ h^{-1}\mathrm{Mpc}$ to minimize the residual RSD effects. The choice of $R_{\mathrm{s}}$ is motivated in the Appendix A. As for the number of cells, we verified that its choice has no significant impact on the reconstruction as long as the size of the cell is smaller than the smoothing radius (White 2021). The fact that the smoothing radius is just 1.5 smaller than that of the smallest void in the sample ($\sim 7.5\ h^{-1}\mathrm{Mpc}$) has no significant impact on our results. The role of the smoothing is to minimize the effects of nonlinearities when removing RSDs, which it does effectively, as shown in the following sections. Importantly, it plays no role in the void detection process, which relies on a Voronoi tessellation of the discrete tracer distribution. As in the previous section, we removed all halos displaced outside the cube resulting in a loss of $\sim 0.4\%$ with respect to the original catalog in real space.
@@ 3.4 The void catalogs
Void analyses identify these structures in a 3-dimensional distribution of mass tracers, in our case simulated dark matter halos. Several void identification algorithms have been proposed in the literature, (see e.g. Colberg et al. 2008; Cautun et al. 2014, for a cross-comparison of various techniques available at that time). In this work, we identified voids in the simulated halo catalogs with the publicly available code VIDE  [^4] https://bitbucket.org/cosmicvoids/vide_public/ (Sutter et al. 2015), based on the watershed algorithm ZOBOV (Neyrinck 2008), which has been widely tested and applied in the literature (e.g. Hamaus et al. 2020; Contarini et al. 2023; Verza et al. 2024). This void finder estimates local densities using Voronoi tessellation and identifies voids as basins in the tracer field. The output is a catalog of voids, each described by its center
@@
calculated as the volume-weighted barycenters, and the effective radius
@@
We applied VIDE to all three types of catalogs (real, redshift, and reconstructed space) across 100 mocks. Fig. 1 on the top panel shows the void size function, i.e., the abundance of voids as a function of their effective radius calculated as the number density of voids $n_{\mathrm{v}}$ per unit radius interval, in real (green stars), redshift (orange triangles), and reconstructed (blue dots) space. The solid lines represent the mean across 100 mocks, while the shaded regions indicate the standard deviation. The results in real and reconstructed space are in good agreement within errors, whereas redshift space shows a slight mismatch, as can be better appreciated by looking at the difference with the real space void size function, normalized by the $1\sigma$ error, represented in the bottom panel of Fig. 1. The discrepancies between redshift and real space are particularly prominent at intermediate scales $20,40]\ h^{-1}$Mpc where the voids are more abundant, while residuals between real and reconstructed space do not show a dependence on the size, always being within the $1\sigma$ interval (grey band). This suggests that the reconstruction procedure successfully recovers the statistical properties of the void distribution, including both their abundance and size, while redshift space distortions affect their identification. A possible explanation is that voids squished along the LOS in real space are more likely to not be selected in real space while identified in redshift space, as discussed in Correa et al. (2022). Additionally, void finders identify spurious, shallow voids due to shot noise from the discrete tracer distribution, commonly referred to as Poisson voids (Neyrinck 2008; Cousinou et al. 2019). These will impact void statistics such as the void size function and the VGCF. A standard mitigation is to exclude small voids below a fixed multiple of the mean tracer separation (mps) as in Hamaus et al. (2020), with values typically around 3 times the mps. However, such cuts are somewhat arbitrary and not fully effective at eliminating contaminants, while considerably reducing the sample size (Cousinou et al. 2019). For this reason, we chose to retain all identified voids and test the robustness of our analysis by comparing results with and without small voids (see Section 5.2). A more detailed investigation into the impact of spurious voids is left for future work.
@@ 4.1 Estimator
The multipoles of the VGCF are obtained using the Davis-Peebles (DP) auto-correlation function estimator (Davis and Peebles 1983). The DP estimator is based on counting object pairs at a given separation and comparing them with pair counts performed in a population of synthetic objects distributed over the same volume and sharing the same selection effects as the real ones but with no intrinsic spatial correlation properties, the so-called random sample.
@@
where $\mathcal{D}_{\text{v}}\mathcal{D}_{\mathrm{h}}$ are the void-halo (or more generally void-tracer) pair counts at radial $r$ and angular $\mu$ separation, and $\mathcal{D}_{\text{v}}\mathcal{R}_{\mathrm{h}}$ are the void-random pair counts. The quantities $n_{\mathrm{R}}$ and $n_{\mathrm{h}}$ represent the total number counts of the random objects and of halos respectively. In Eq. (22) $r$ is a dimensionless quantity representing the physical separation between the center of the void and the halo divided by the void effective radius $R$. In this analysis the pair counts are evaluated in 25 equally-spaced bins in the range $[0,3]$ while $\mu$ is evaluated in $100$ bins between $[0,1]$. In our analysis, we computed the cross pair with the random catalog by generating a single catalog of synthetic objects uniformly distributed across the volume, with a number density 20 times higher than that of the halos. This random catalog is used for all cross-correlation analyses, as they share the same geometry and mean halo density.
Once the VGCF is evaluated, the multipoles can be computed using Eq. (16). Particular attention is given to the quadrupole, as it serves as a crucial diagnostic tool for understanding the effects of RSD. In real space, where no RSD are present, the quadrupole signal is expected to vanish due to the spherical symmetry of the stacked void profile. Thus, a zero quadrupole signal is used as a benchmark to assess the accuracy of the reconstruction procedure within the appropriate cosmological model. In addition to the impact of RSD on the quadrupole shape, AP geometric distortions cause the quadrupole to deviate from zero. The direction of the deviation, compression or elongation, depends on whether the parameter $\varepsilon$, defined by Eq. (4), is lower or greater than 1. As we use the fiducial cosmology of the simulation in the distance estimate, no genuine AP-induced distortion is expected, and $\varepsilon$ should therefore be unity.
@@ 4.2 Likelihood
To sample the posterior probability distribution of all model parameters we make use of the affine-invariant Markov chain Monte Carlo (MCMC) ensemble sampler emcee (Foreman-Mackey et al. 2019). The quality of the maximum-likelihood model (best fit) is assessed via evaluation of the reduced $\chi^{2}$ statistics
@@ 4.3 Covariance matrix
All analyses are carried out using a dedicated covariance matrix for each mock, estimated via jackknife resampling of its corresponding void catalog. Given that the number of available mocks is insufficient to reliably compute a full covariance matrix across realizations, we followed a standard approach adopted in the literature and apply a delete-1 jackknife strategy to estimate the covariance for each of the 100 mocks individually (see e.g. Hamaus et al. 2020, 2022; Radinović et al. 2023). In Appendix A.2, we show that the jackknife-derived covariance matrices are consistent with those obtained from the full mock-to-mock covariance. Based on this agreement, we adopt the jackknife covariance, $\text{Cov}[\boldsymbol{\xi}]_{ij}$, throughout our analysis.
This methodology relies on ergodicity, which allows us to average measurements across different spatial patches to estimate the covariance matrix (Hamaus et al. 2020). In this approach, we remove one void at a time from the sample when estimating $\xi_{\ell}(r)$, providing a total of $N_{\mathrm{v}}$ jackknife samples, where $N_{\mathrm{v}}$ is the total number of voids in the catalog. These samples can then be used in Eq. (27) to calculate $\text{Cov}[\boldsymbol{\xi}]_{ij}$, with an additional normalization factor of $(N_{\mathrm{v}}-1)$ to account for the statistical weight of the jackknife sample size. The elements of the covariance matrix are computed as follows
@@ 5 Results
  • The first – a pre-requisite for the second – is to assess the performance of the reconstruction algorithm in removing RSDs. To do so, we focus on the parameter $\beta$, which quantifies the amplitude of RSD effects in the VGCF multipoles. In the ideal case of perfect RSD removal, we expect $\beta=0$. In practice, however, uncertainties are expected to arise from imperfect dynamical reconstruction, void detection, and VGCF estimation. To isolate the effects introduced solely by reconstruction and assess their impact on the void-finding procedure, we compared the best-fit values of $\beta$ obtained from the reconstructed catalogs with those measured directly in real space, rather than with the expected value $\beta=0$. A close match between the two $\beta$ values would indicate that the reconstruction effectively removes RSDs.
  • The second is to evaluate whether performing the AP test in reconstructed space yields tighter constraints on the geometric distortion parameter $\varepsilon$, compared to using voids identified in redshift space. To this end, we compared the AP constraints derived from both redshift and reconstructed spaces void catalogs using the full void sample (Section 5.1). We then assessed the impact of excluding small voids, motivated by their greater susceptibility to noise and systematics, by repeating the analysis on size-selected subsamples in both spaces (Section 5.2).
Finally, to assess the robustness of the results presented in this section, we performed a series of consistency tests. These include examining the sensitivity of the method to the parameters used in the reconstruction process and evaluating its stability with respect to the choice of the covariance matrix. A detailed description of these tests and their outcomes is provided in Appendix A.
@@
Fig. 2, shows the three measured multipoles averaged for the 100 mocks (described in Section 3) – in redshift space (orange triangles and solid curves), reconstructed space (blue dots and solid curves) and real space (green stars and solid curves). Focusing on the quadrupole (central panel), we observe a positive bump near the ridge at $r\sim 1$ in redshift space. This feature signals the presence of coherent outflows and manifests as an apparent elongation of the void along the LOS. On the other hand, the quadrupole measured in reconstructed space is compatible within error bars with the one in real space, with both being consistent with zero, indicating the effective removal of RSDs. The consistency of the hexadecapole with zero further supports this conclusion.
@@
One of the central motivations of this work is to develop a methodology that enables the estimation of the geometric distortion parameter $\varepsilon$ independently of the systematics that affect the measurement of $\beta$ in redshift space, systematics that can drive $\beta$ far from its true value. Disentangling these contributions is particularly important because, in redshift space, the parameters $\varepsilon$ and $\beta$ are strongly correlated. This is evident in Fig. 3, where their joint posterior displays a pronounced alignment, with a correlation coefficient of $\mathcal{R}=0.67$ estimated directly from the parameter chains. Since performing a robust AP test requires an accurate treatment of RSD, applying reconstruction methods proves highly effective: as shown in Fig. 4, the posterior of $\varepsilon$ and $\beta$ becomes nearly spherical after reconstruction – and the correlation is reduced to $\mathcal{R}=0.23$ – demonstrating that the degeneracy between RSD and AP effects is largely mitigated.
@@
Table 1: Parameter constraints for $\varepsilon$ and $\beta$. [^5] Table summarizing the values of the parameters $\varepsilon$ and $\beta$, mean values with $68\%$ confidence intervals, for the analyses in redshift, reconstructed, and real space. The precision in reconstructed space is increased by $\sim 23\%$ with respect to redshift space, and approaches the real-space analysis precision.